OpenStudy (anonymous):

5. The graph of g(x) results when the graph of f(x) = |x| is shifted 4 units to the right. i got b Which is the equation of g(x)? (Points : 4) g(x) = 4|x| g(x) = |x| + 4 g(x) = |x 4| g(x) = |x + 4|

OpenStudy (jdoe0001):

y = A ( Bx + C ) + D C = horizontal shift, C > 0, to the left, C < 0, to the right

OpenStudy (anonymous):

i dont understand what you just said, is this suposed to be in y = mx + b? or f(x)

OpenStudy (jdoe0001):

hmmm ok, then ... y = A | Bx + C | + D C = horizontal shift, C > 0, to the left, C < 0, to the right

OpenStudy (jdoe0001):

that's what the translation, transformation or shifts so-called, occur

OpenStudy (jdoe0001):

you need it shifted 4 units to the right.... so C < 0, or C is negative

OpenStudy (anonymous):

so you have l x + 4l or something like that?

OpenStudy (jdoe0001):

|x+4| will be a shift to the left

OpenStudy (anonymous):

- would be the shift to the right

OpenStudy (jdoe0001):

so the obvious now will be | x - 4 |

OpenStudy (jdoe0001):

yeap

OpenStudy (anonymous):

so the answeri would have would be -lx-4l

OpenStudy (jdoe0001):

hmm well... if you just need | x | shifted to the right 4 units, it'd be just | x - 4 |

OpenStudy (anonymous):

sorry the negative did not show up

OpenStudy (jdoe0001):

-|x - 4 | will shift it 4 units to the right, and will also flip it upside down

OpenStudy (anonymous):

5. The graph of g(x) results when the graph of f(x) = |x| is shifted 4 units to the right. i got b Which is the equation of g(x)? (Points : 4) g(x) = -4|x| g(x) = -|x| + 4 g(x) = -|x- 4| g(x) = -|x + 4|

OpenStudy (jdoe0001):

well.... "b" will be a vertical shift

OpenStudy (jdoe0001):

-|x| + 4 is |x| flipped upside down and 4 units UP

OpenStudy (jdoe0001):

unless f(x) = -|x|

OpenStudy (anonymous):

it does f(x) = -lxl sorry

OpenStudy (jdoe0001):

then now you know which one it's :)

OpenStudy (anonymous):

-lxl + 4?

OpenStudy (jdoe0001):

recall that y = A | Bx + C | + D C = horizontal shift, C > 0, to the left, C < 0, to the right

OpenStudy (anonymous):

wait if -|x - 4 | will shift it 4 units to the right,will it still flip it upside down?

OpenStudy (jdoe0001):

well, it'll shift it to the right 4 units, it won't flip it upside down, not because original it was -|x|

OpenStudy (anonymous):

oh ok i see now i tried to graph this some how and thats why i got b, thank you again :)

OpenStudy (jdoe0001):

yw